OpenStudy (anonymous):

Find the angle with a tangent ratio of 0.4877.

OpenStudy (e.mccormick):

OK, that is positive, so what quadrants have positive tangents?

OpenStudy (e.mccormick):

Once you know that, then: \(\tan\theta =0.4877\implies\) \(\arctan(\tan\theta) =\arctan(0.4877)\implies\) \(\theta =\arctan(0.4877)\) After that, you put the two facts together to find the possible angles that answer your question.